OpenStudy (anonymous):

Let {a_n} and {b_n}be sequences of positive real numbers such that for n≥2, a_n+1=1/2(a_n +b_n), b_n+1=sqrt(a_n x b_n) (a) Show that for 2n≥, {a_n} is monotonic decreasing and {b_n}is monotonic increasing. (b) Deduce that {a_n} and {b_n} have the same limit. (Hint: Refer to the Arithmetic-Geometric Inequality to get a relation between a_n and b_n)

OpenStudy (anonymous):

wiah... tatau kene la tanye kalau tau xtaye la

OpenStudy (anonymous):

wiahhhh i nak berjoli sat i tau i balik i nk u sdia kan jwpn hahahahah

OpenStudy (anonymous):

\[a _{n+1}=1/2(a _{n}+b _{n})\] \[b _{n+1}=\sqrt{a _{n}b _{n}}\] is this the sequence?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[a_{n+1}=\frac{1}{2}(a_n +b_n), b_{n+1}=\sqrt{a_n x b_n}\]like this?

OpenStudy (anonymous):

like myko equation

OpenStudy (anonymous):

\[a _{n+1}-a _{n}=1/2(a _{n}+b _{n})-a _{n}=b _{n}/2-a _{n}/2\] so for the 1º question you have to prove that this is negative

OpenStudy (anonymous):

i guess we start with \[\frac{a_n+b_n}{2}\geq \sqrt{a_nb_n}\]

OpenStudy (anonymous):

\[b _{n+1}^{2}/b _{n}=a _{n}\] substituting in the one befor: \[b _{n}-b _{n+1}^{2}/b _{n}=(b _{n}^{2}-b _{n+1}^{2})/b _{n}\]

OpenStudy (anonymous):

so looks if bn is increasing, an is decreasing

OpenStudy (anonymous):

but also oposite is true....

OpenStudy (anonymous):

maybe @satellite73 has any ideas....

OpenStudy (anonymous):

i tried this we know that \[\frac{a_n+b_n}{2}\geq \sqrt{a_nb_n}\] as arthithmetic mean is greater then geometric one or at least we know that \[\frac{a_1+b_1}{2}\geq \sqrt{a_1b_1}\] and therefore \[a_2\geq b_2\]

OpenStudy (anonymous):

by induction (easy induction) you can prove that therefore \[a_n\geq_n\]

OpenStudy (anonymous):

sorry i meant \[a_n\geq b_n\]

OpenStudy (anonymous):

but what about increasing, decreasing

OpenStudy (anonymous):

now \[a_{n+1}=\frac{a_n+b_n}{2}\] and therefore \[2a_{n+1}=a_n+b_n\] \[2a_{n+1}\leq 2a_n\] \[a_{n+1}\leq a_n\]

OpenStudy (anonymous):

inequality since \[b_n\leq a_n\]

OpenStudy (anonymous):

oh wait i did i get that backwards? i have to run, but look it over

OpenStudy (anonymous):

no no it is right

OpenStudy (anonymous):

inequality is correct, you can check it

OpenStudy (anonymous):

and how to deduce the sequences hve same limit??

OpenStudy (anonymous):

If we agree that \[ a_n\] and \[ b_n\] have limits a and b then both sides \[ a_{n+1} = \frac {a_n+b_n}2 \] will converge to \[ a =\frac { a+b}2 \] which implies that a=b

OpenStudy (anonymous):

This sequence converges very fast. Example \[ b_1=1, a_1=1000\] Here are the first 10 terms of the sequence a \[ \left\{1000,\frac{1001}{2},266 .061388300841896660,195.933 793615277881874,\\189.4439047 7151416056,\\189.388306490415 24224,189.38830240995089754 ,189.38830240995087556,\\189. 38830240995087556,189.38830 240995087556\right\} \] Here are the first 10 terms of the sequence b \[ \{1,31.622776601683793320,125. 80619892971386709,\\ 182.95401 592775043925,189.3327082093 1632393,189.388298329486552 83,\\ 189.38830240995085358,18 9.38830240995087556,\\189.388 30240995087556,189.38830240 995087556\} \]

OpenStudy (anonymous):

ya..i can see it..thanks..

OpenStudy (anonymous):

how we know that \[a_{n+1}\le a_{n}\] only for \[n \ge2\]?

OpenStudy (anonymous):

Notice that \[ b_{n+1} =\sqrt{ a_n b_n} \le \frac {a_n+b_n}{2}= a_{n+1}\\ a_{n+2} =\frac {a_{n+1}+b_{n+1}}{2}\\ so\\ b_{n+1} \le a_{n+2} \le a_{n+1} \] by being the mid point of the segment \[ [b_{n+1} ,\, a_{n+1} ] \]

OpenStudy (anonymous):

wow still on this one good

OpenStudy (anonymous):

i notice that we did not show that \(b_n\) is decreasing, nor did we show that there is a limit, although if you have an increasing sequence bounded above then it must have a limit, likewise a decreasing sequence bounded below

OpenStudy (anonymous):

sorry \(b_n\) increasing

OpenStudy (anonymous):

i still dont undrstnd how we know \[a_{n} \] is monotonic decreasing for \[n \ge2\]

OpenStudy (anonymous):

ya..then?

OpenStudy (anonymous):

we kn ow \[\frac{a+b}{2}\geq \sqrt{ab}\]

OpenStudy (anonymous):

so \[a_2=\frac{a_1+b_1}{2}\geq b_2=\sqrt{a_1b_1}\] yes

OpenStudy (anonymous):

more generally \[a_{n+1}=\frac{a_n+b_n}{2}\geq b_{n+1}=\sqrt{a_nb_n}\]

OpenStudy (anonymous):

in other words it is always true that \(a_n\geq b_n\)

OpenStudy (anonymous):

we know \[a_{n+1}=\frac{a_n+b_n}{2}\] and therefore \[2a_{n+1}=a_n+b_n\]

OpenStudy (anonymous):

now since \(b_n<a_n\) if we replace \(b_n\) by \(a_n\) in the above equality \(2a_{n+1}=a_n+b_n\) we get the inequality \[2a_{n+1}\leq a_n+a_n\]

OpenStudy (anonymous):

dividing by 2 we get \[a_{n+1}\leq a_n\] which is what we wanted

OpenStudy (anonymous):

how can we say that it is always true for \[a _{n}\ge\]\[b _{n}\]??

OpenStudy (anonymous):

i meant \[a _{n}\ge b _{n}\]

OpenStudy (anonymous):

is it true if we say \[a_{n}\] is monotonic decreasing for \[n \ge1\]??

OpenStudy (anonymous):

yup. how can we ensure all this is about n≥2? how it is related actually?

OpenStudy (anonymous):

is it based on induction??

OpenStudy (anonymous):

still need induction? i'm very weak in the induction part.